|IODP Proceedings Volume contents Search|
|Expedition reports Research results Supplementary material Drilling maps Expedition bibliography|
The science portion of Expedition 311 officially began with the first line ashore in Astoria, Oregon (USA), at 1306 h on 15 September 2005. After completing port call activities, we departed Astoria at 0944 h on 19 September. The 188 nmi transit to Site U1325 was completed in 19.5 h at an average speed of 9.6 kt. Our initially planned drilling location was scheduled to be proposed Site CAS-03B (Collett et al., 2005), but a large slump identified from newly available multibeam data led us to switch priority to alternate Site U1326 (proposed Site CAS-03C). Approval depth (from IODP's Environmental Protection and Safety Panel and territorial authorities) at Site U1326 was 300 mbsf, 50 m shallower than our preferred penetration to ensure that the entire LWD/MWD tool string penetrated through the interval of interest. Therefore, we decided to change the drilling sequence, occupying Site U1325 prior to Site U1326, to allow more time for approval of the late depth-extension request, which did not arrive before completion of the hole.
The LWD/MWD tool string was assembled upon arrival on site, and the drill string was run to the seafloor. During function testing of the tool string, a leak in the top drive swivel wash pipe packing was discovered, which required 1.5 h to repair. Hole U1325A was spudded at 2310 h on 20 September 2005, at a seafloor depth of 2212.0 meters below rig floor (mbrf). Initial spud-in parameters were set at 290 gallons per minute (gpm) circulation and 10–20 rpm, which was increased to 40 rpm at 10 mbsf, and then again to 60 rpm at 29 mbsf to maintain a 25–30 m/h rate of penetration (ROP) while also maintaining a minimum of 290 gpm circulation. The hole was logged to 350 mbsf without incident and within the prescribed protocol guidelines, requiring no corrective actions (see "Downhole logging" in the "Methods" chapter). After displacing the hole with 10.5 ppg sepiolite mud, the drill string was tripped to the rig floor so that the digital log data from the LWD/MWD logging tools could be downloaded while transiting in dynamic positioning mode to Site U1326. The bit cleared the rotary table at 0515 h on 22 September, officially ending operations in Hole U1325A.
After pulling the drill pipe back to 1000 mbrf, the ship was moved 5.5 nmi in dynamic positioning mode from Site U1328 to Site U1325. The arrival time was 1430 h on 17 October 2005 (Fig. F6). Prior to starting the hole, a water-sampling temperature probe (WSTP) water sample was collected. Hole U1325B was spudded with the advanced piston corer (APC) system at 1910 h. The estimated water depth was 2194.9 meters below sea level (mbsl; 2206.2 mbrf), although the core was almost a full core (see "Hole U1325D") (Table T1). When attempting to recover Core 2H, the inner core barrel became stuck near the drill bit. It took several hours to work the core barrel loose. The core barrel was returned slightly bent with a kinked wireline. The second APC core contained a relatively large amount of fine sand, which probably caused the barrel to rebound. As a result, the top of the APC assembly became entangled with the wireline. After replacing part of the barrel assembly and reterminating the coring wireline, an XCB core barrel was deployed to work the hole below the sand section identified in the second APC core.
The hole was advanced by combination of APC, XCB, and pressure coring systems to a depth of 206.5 mbsf. Three deployments of the third-generation advanced piston corer temperature (APCT-3) tool returned very high quality data (Table T1). At 71.5 mbsf, we had an incomplete stroke of the APC system for Core 8H and switched to XCB coring. A Davis-Villinger Temperature Probe (DVTP) run at 140.5 mbsf yielded high-quality temperature data. Recovery dropped dramatically in the last three XCB cores, averaging 6% (Table T1). During Hole U1325B operations at 1230 h on 18 October, marine mammal observer, Scott Toews, departed the ship by helicopter.
Five pressure coring tools were deployed in Hole U1325B. The first PCS coring run was taken at 82.4 mbsf but did not return with a pressurized core. The HRC was deployed at 129.9 and 197.4 mbsf, but both runs failed to return core under pressure. We also deployed the FPC at 169.4 mbsf, but it did not retrieve a pressurized core. The last PCS deployment was at 206.1 mbsf, but the tool became jammed in the bottom-hole assembly (BHA) after penetrating just 40 cm into the formation. Hole U1325B was terminated after four unsuccessful tool recovery attempts. Operations in Hole U1325B officially ended at 2125 h on 19 October when pipe was tripped back to the seafloor. Upon inspection, the PCS bit appeared to have overheated, which caused some melting and enlarged the bottom of the barrel enough to prevent it from being pulled back in the pipe.
After checking for potential problems with the BHA, the drill string was tripped to the seafloor, and Hole U1325C was spudded at 0820 h on 20 October 2005. The hole was drilled to 188 mbsf, where core recovery in Hole U1325B began to deteriorate. Coring operations resumed with the deployment of the XCB system at 2115 h, and the first core contained only 3.28 m of sediment. We decided to try cutting XCB half-cores to see if this would increase core recovery. The first short XCB core recovered only 0.73 m of sediment, and the second one recovered none. Thereafter, we decided to switch back to standard full-length XCB core deployments. XCB coring operations interspersed with two pressure coring runs and two DVTP deployments deepened the hole to a total depth (TD) of 304.3 mbsf. The last core was brought on deck at 0715 h on 22 October. An FPC pressure core was deployed at a depth of 217.6 mbsf, but it did not return with sediment under pressure. A PCS pressure core was successfully recovered from 256.1 mbsf. The DVTP tool was deployed at 218.6 and 275.6 mbsf.
To prepare for logging, a mud sweep and wiper trip were conducted. The hole was displaced with weighted mud, and the pipe was pulled to logging depth (76.2 mbsf). During the previous evening and early morning hours of 22 October, sea state severely deteriorated and resulted in a sustained ship heave of 5 m. After reviewing the weather forecast and discussing options, we decided that two separate logging runs would be attempted with tools without calipers to reduce the potential risk of tool string damage. The first run included the Phasor Dual Induction Tool (DIT) and the Hostile Environment Gamma Ray Sonde (HNGS). The second run included the Dipole Sonic Imager (DSI), the Scintillation Gamma Ray Tool (SGT), and the Temperature/Acceleration/Pressure (TAP) tool. The first tool string was run in the hole at 1325 h, but we were only able to lower it to 259.8 mbsf, 44.5 m shallower than total coring depth. The run was successfully completed despite some difficulty reentering the drill pipe past the flapper valve, and the first tool string was back on deck at 1800 h. The modified sonic tool string (without the FMS) was deployed at 2000 h, but it reached a TD of only 185.8 mbsf. The available hole was logged successfully. The repeat pass of the sonic tool string reached 183.0 mbsf and the tool was back to the rig floor at 0030 h. After rigging down logging, the bit was pulled clear of the seafloor at 0200 h on 23 October, ending operations in Hole U1325C.
To ensure that a mudline core was obtained at Site U1325, one APC core was taken with the bit at 2200.0 mbrf. Hole U1325D was spudded at 0335 h on 23 October 2005. Core 1H recovered 4.69 m and indicated a seafloor depth of 2193.2 mbsl (2204.8 mbrf). This depth suggested that Hole U1325B was initiated ~1.4 m below the seafloor. The heave state throughout Site U1325 operations, however, was 3 m or higher, increasing the uncertainty of seafloor depth calculations using the recovery of the mudline core. Shipboard analyses of Core 1H indicated that Hole U1325B may have missed the mudline by as much as ~2.5 m (see "Interstitial water geochemistry"). The drill pipe was pulled clear of the seafloor at 0430 h on 23 October, ending operations at Site U1325.